# The SHMS 11 GeV/c Spectrometer in Hall C at Jefferson Lab

### v1

@@ 4.7.2 Construction
The SHMS calorimeter consists of two parts (see Fig. 39): a Preshower at the front of the calorimeter, and the main part, the Shower, at the rear. An expedient and cost-effective choice was to use modules from the decommissioned HERMES calorimeter for the Shower part, and modules from the decommissioned SOS calorimeter for the Preshower. With these choices, the Preshower is 3.6 radiation lengths thick, and the Shower is 18.2 radiation lengths deep. The combination almost entirely absorbs showers from 10 GeV electrons.
@@ 4.7.3 Choice and studies of PMT bases
The Preshower PMT high voltage base is optimized for good linearity (better than 1%), high rate capability and a weak variation of PMT gain with anode current [59]. A design was selected which is a purely resistive, high current (2.3 mA at 1.5 kV), surface mounted divider (), operating at negative HV. The relative applied HVs down the dynode chain (from cathode to anode) are: 3.12/1.50/1.25/1.25/1.50/1.75/2.00/2.75/2.75. The supply voltage for a gain of is approximately 1750 V.
@@ 4.7.4 Monte Carlo simulations
Prior to construction, the calorimeter design was simulated in order to optimize the setup and get predictions for key characteristics. The simulations were based on the GEANT4 package [60], release 9.2. As in the simulations of the HMS calorimeter (see [58]), the QGSP_BERT physics list was chosen to model hadron interactions [61]. The code closely followed the parameters of the detector components. Other features are added into the model to make it more realistic, such as:
@@
The Cherenkov light propagation and detection were handled by a custom code, using an approximation of strict rectangular geometry of the lead glass blocks with perfectly polished surfaces. Light reflection and absorption by the Mylar wrapping was modeled via aluminum complex refractive index, with Mylar support facing the block, and a thin air gap between the wrapping and the block. Both light passage to the PMT photocathode through the optical grease and the PMT window, and reflections from the block sides, were modeled using the approximation of thin dielectric layers ([62], p. 360). The electronic effects, such as pedestal widths and channel to channel PMT gain variations, were assumed as for the HMS calorimeter before the 12 GeV modifications.
@@ 4.7.5 Calorimeter Gain Matching
As gain matching had to be achieved by adjustment of high voltages on the PMT bases, knowledge of gain versus supplied HV was needed. These were obtained by measuring signals from MIP pions at two supply high voltages for all the Shower channels, at 1.4 kV and 1.5 kV (see Fig. 41). By assuming gain dependence on the supplied voltage of proportional to [51], the average exponent was found to be 5.70 0.01 for a subset of 100 channels.
@@ 4.7.6 Calorimeter Calibration
The calorimeter calibration procedure corrects for the gain differences between channels. Good electron events are selected by utilising the gas Cherenkov detector(s). The standard calibration algorithm [63] is based on minimization of the variance of the estimated energy with respect to the calibration constants, subject to the constraint that the estimate is unbiased (relative to the primary energy).
@@
is referred to as the normalized energy. For electrons, should be close to 1. An example of the normalized energy distribution for electron tracks can be seen before and after a successful calibration in Figs. 45 and 46.
@@ 5 Trigger and Data Acquisition
The Hall C data acquisition (DAQ) system is designed to meet the needs of a high luminosity, dual spectrometer (SHMS + HMS) configuration, with the capability of extracting polarization-dependent absolute cross sections with precision at the 1% level or better. JLab’s CODA data acquisition software [64] provides a framework that ties together a distributed network of read-out controllers (ROCs) controlling multiple crates of digitization hardware, event builders to serialize the data, and event recorder processes to write the data to disk. It also provides a graphical control interface for the users.
The Hall C DAQ system can run in dual-arm trigger mode that requires a coincidence between both spectrometers, or each arm’s DAQ may be run entirely independently. Incorporating additional detector systems into the standard two-arm design is also straightforward. A high-level block diagram of trigger formation and readout for each spectrometer arm (SHMS or HMS) is depicted in Fig. 47.
@@ 5.1 Standard Triggers
All trigger-related PMT signals from both the SHMS and HMS are routed out of the experimental Hall to a dedicated electronics room on the main level of the Hall C Counting House using low-loss RG-8 air-core signal cables. Those signals are then split with one output running into a JLab F250 flash analog to digital converter (FADC)[53], while the second output is processed and discriminated. All discriminated pulses are delivered to scalers for rate information, TDCs for precision timing measurement, and to form pre-triggers as described below. This design allows direct access to all raw signals that may participate in a trigger during beam operations and has proven invaluable during the debugging and commissioning phases of Hall operations.
Non-trigger related signals include wire-chamber readouts and the Shower (but not Preshower) layer of the SHMS calorimeter. The readout electronics for those sub-detectors remain inside their respective detector huts within the experimental Hall. All SHMS calorimeter PMT signals are fed into F250 FADCs configured to provide timing, integrated energy, pulse amplitude, and (optionally) pulse profile data as desired. The wire-chamber timing signals are digitized using multi-hit CAEN v1190 modules [65].
@@ 5.1.1 JLab F250 Flash ADCs
The JLab F250 flash ADC modules are an FPGA-based design developed by the Jefferson Lab Fast Electronics group [53] and are used lab wide. Each F250 module provides 16 independent 50 input channels. The voltage at each input channel is continuously digitized into an 8 s ring buffer at 250 MHz, with a resolution of 12 bits, and a hardware adjustable full-scale range. When a module receives a readout trigger, digitized sample data stored in the ring buffer are processed in parallel without incurring front-end deadtime. In typical operation each ‘hit’ over a pre-programmed threshold is assigned an interpolated leading-edge threshold time (1 ns resolution), integrated energy (analogous to a charge-integrating ADC value), a peak amplitude, and a measurement of the DC offset (pedestal) present on the channel prior to the detected pulse. Full pulse-profile data for each hit may also be stored if desired. However, that mode increases the data rate by several orders of magnitude, and is generally used only for debugging or limited duration pulse characterization runs.
@@ 5.1.2 SHMS Triggers
An optional SHMS aerogel Cherenkov detector, as detailed in Sec. 4.6, may also be installed. It employs seven PMTs on each side of its diffusion box. The signals from all 14 PMTs are handled analogously to the gas Cherenkov detectors, with each analog signal being split and read out by an individual FADC channel, with the second outputs being summed and discriminated to form an associated aerogel pre-trigger. The pre-trigger is routed to a scaler and v1190 TDC as well.
A block diagram for the Cherenkov pre-triggers is presented in Fig. 50.
The SHMS Preshower layer, described in Sec. 4.7, consists of 28 lead-glass blocks arranged in 14 rows, with 2 blocks to a row. Each block is coupled to a single PMT on the side facing the perimeter of the layer. Analog signals from the 28 PMTs are split and summed in 3 groups of 4 rows, and 1 group of 2 rows. Each of the 4 group sums is read out by an FADC channel for cross checks. The 4 group sums are summed in turn to provide a total Preshower sum which is then discriminated and provides the SHMS PSh pre-trigger. Provision is made to generate independent pre-triggers for both low and high energy depositions in the Preshower layer (PSh_Lo and PSh_Hi, respectively) as seen in Fig. 51.
@@ 5.3 Auxiliary Data Collection
The standard method for slow controls data logging is through the Experimental Physics and Industrial Control System (EPICS) [66]. EPICS is a system of open source software tools and applications used to provide control user interfaces and data logging for systems such as high- and low-voltage detector power supplies, target systems, spectrometer magnets, vacuum, and cryogenic systems, etc.
@@ 5.4 Online Hall C Computing Environment
All systems have direct access to the JLab centrally managed Scientific Computing resources. This includes multi-petabyte tape storage and online disk facilities, as well as a several thousand core compute farm for simulation and offline data analysis [67].
@@ 6 Software
Hall C Data is analyzed by the Hall C analysis package hcana. This package does full event reconstruction for the SHMS used alone or in coincidence with other detectors. hcana is based on the modular Hall A analyzer [68] ROOT [69] based C++ analysis framework. This framework provides for run time user configuration of histograms, ROOT tree contents, cuts, parameters and detector layout.
@@ 6.1 Online Monitoring
After each data taking run is started, a subset of the data is analyzed with hcana. An easily configurable histogram display GUI is used to view diagnostic histograms and compare them to reference histograms. The EPICS [66] control system alarm handler is used to monitor experiment settings and beam conditions. This includes spectrometer magnet settings, detector high voltages, drift chamber gas, cryogenic systems and spectrometer vacuum.
@@ 7.3.1 Hodoscope Performance
Beam data confirmed the expectation that the detection efficiency for low momentum protons, for example, will be smaller than that for pions or electrons simply due to the reduced number of Cherenkov photons that particles close to their firing threshold will produce. This is exemplified by Figs. 61, 62 and 63.
@@ 7.3.2 Drift Chamber Performance
The SHMS drift chambers have proven to provide reliable tracking for electrons and hadrons across a broad range of momenta. The position resolution depends on how well the wire planes are aligned in the reconstruction software, but per-plane resolutions of 175 m are typical. The drift chambers have also performed very well at high rate, with tracking efficiencies exceeding 96%, even at pre-trigger rates over 2 MHz. The tracking efficiency as a function of the S1X hodoscope trigger rate (a good proxy for the overall event rate) in the SHMS can be seen in Figs. 64 and 65.
@@ 7.3.3 HGC Performance
Lastly, measurements of the efficiency across a variety of momentum settings can be used to verify the index of refraction of the Cherenkov media. The relationship between efficiency and momentum is fit with the equation [71]
where is the detector efficiency, is the momentum of the , and and are free parameters. Data taken in the range of 2.53 GeV/c to 5.05 GeV/c with the HGC filled with C4F10 yields an index of refraction of . This is in agreement with the accepted value of [72]. Additional performance details are given in [73].
@@ 8 Conclusion
Numerous experiments have completed and published high profile results since the SHMS was commissioned in 2017. This includes many high profile results on color transparency [74, 75], the EMC effect [76], deuteron structure [17] and proton structure [77]. This also extends to detailed studies of the proton’s gravitational form factors [78]. Results on the F2 ratio for D/H[79] and the flavor dependence in charged pion SIDIS[80] were recently submitted for publication.
Many more high profile scientific results are expected in the near future, with several experimental campaigns now completed and data analysis in advanced stages. Due to the design parameters of the SHMS, it could also be utilised extensively in an upgraded, 20+ GeV Jefferson Lab scenario. Some possible experiments that utilise the SHMS in a 20+ GeV era are outlined in the 22 GeV white paper [81].

### v2

@@ 4.7.1 Design and Construction
The SHMS calorimeter consists of two parts (see Fig. 40): a Preshower at the front of the calorimeter, and the main part, the Shower, at the rear. An expedient and cost-effective choice was to use modules from the decommissioned HERMES calorimeter [Avakian:1998bz] for the Shower part, and modules from the decommissioned SOS calorimeter [mkrtchyan] for the Preshower. With these choices, the Preshower is 3.6 radiation lengths thick, and the Shower is 18.2 radiation lengths deep. The combination almost entirely absorbs showers from 10 GeV electrons. The SHMS Preshower consists of one layer of 28 TF-1 type lead glass blocks stacked in two columns in an aluminum enclosure (not shown in Fig. 40). Each Preshower block is an block of TF-1 lead-glass with a Photonis XP3462B PMT attached. The Shower consists of 224 TF-101 type lead glass modules stacked in a “fly’s eye” configuration of 14 columns and 16 rows. Each Shower block is has a size of and is optically isolated with a Photonis XP3461 PMT attached.
@@ 4.7.2 Choice and studies of PMT bases
The Preshower PMT high voltage base is optimized for good linearity (better than 1%), high rate capability and a weak variation of PMT gain with anode current [Amatuni96]. A design was selected which is a purely resistive, high current (2.3 mA at 1.5 kV), surface mounted divider (), operating at negative HV. The relative applied HVs down the dynode chain (from cathode to anode) are: 3.12/1.50/1.25/1.25/1.50/1.75/2.00/2.75/2.75. The supply voltage for a gain of is approximately 1750 V.
@@ 4.7.3 Monte Carlo simulations
Prior to construction, the calorimeter design was simulated in order to optimize the setup and get predictions for key characteristics. The simulations were based on the GEANT4 package [geant4], release 9.2. As in the simulations of the HMS calorimeter (see [mkrtchyan]), the QGSP_BERT physics list was chosen to model hadron interactions [qgsp-bert]. The code closely followed the parameters of the detector components. Other features are added into the model to make it more realistic, such as:
@@
The Cherenkov light propagation and detection were handled by a custom code, using an approximation of strict rectangular geometry of the lead glass blocks with perfectly polished surfaces. Light reflection and absorption by the Mylar wrapping was modeled via aluminum complex refractive index, with Mylar support facing the block, and a thin air gap between the wrapping and the block. Both light passage to the PMT photocathode through the optical grease and the PMT window, and reflections from the block sides, were modeled using the approximation of thin dielectric layers ([BornWolf], p. 360). The electronic effects, such as pedestal widths and channel to channel PMT gain variations, were assumed as for the HMS calorimeter before the 12 GeV modifications.
@@ 4.7.4 Calorimeter Gain Matching
As gain matching had to be achieved by adjustment of high voltages on the PMT bases, knowledge of gain versus supplied HV was needed. These were obtained by measuring signals from MIP pions at two supply high voltages for all the Shower channels, at 1.4 kV and 1.5 kV (see Fig. 42). By assuming gain dependence on the supplied voltage of proportional to [hamamatsu], the average exponent was found to be 5.70 0.01 for a subset of 100 channels.
@@ 4.7.5 Calorimeter Calibration
The calorimeter calibration procedure corrects for the gain differences between channels. Good electron events are selected by utilising the gas Cherenkov detector(s). The standard calibration algorithm [amatuni] is based on minimization of the variance of the estimated energy with respect to the calibration constants, subject to the constraint that the estimate is unbiased (relative to the primary energy).
@@
is referred to as the normalized energy. For electrons, should be close to 1. An example of the normalized energy distribution for electron tracks can be seen before and after a successful calibration in Figs. 46 and 47.
@@ 5 Trigger and Data Acquisition
The Hall C data acquisition (DAQ) system is designed to meet the needs of a high luminosity, dual spectrometer (SHMS + HMS) configuration, with the capability of extracting polarization-dependent absolute cross sections with precision at the 1% level or better. JLab’s CODA data acquisition software [CODA] provides a framework that ties together a distributed network of read-out controllers (ROCs) controlling multiple crates of digitization hardware, event builders to serialize the data, and event recorder processes to write the data to disk. It also provides a graphical control interface for the users.
The Hall C DAQ system can run in dual-arm trigger mode that requires a coincidence between both spectrometers, or each arm’s DAQ may be run entirely independently. Incorporating additional detector systems into the standard two-arm design is also straightforward. A high-level block diagram of trigger formation and readout for SHMS is depicted in Fig. 48.
@@ 5.1 Standard Triggers
All trigger-related PMT signals from both the SHMS and HMS are routed out of the experimental Hall to a dedicated electronics room on the main level of the Hall C Counting House using low-loss RG-8 air-core signal cables. Those signals are then split with one output running into a JLab F250 flash analog to digital converter (FADC)[JLABF250], while the second output is processed and discriminated. All discriminated pulses are delivered to scalers for rate information, TDCs for precision timing measurement, and to form pre-triggers as described below. This design allows direct access to all raw signals that may participate in a trigger during beam operations and has proven invaluable during the debugging and commissioning phases of Hall operations.
Non-trigger related signals include wire-chamber readouts and the Shower (but not Preshower) layer of the SHMS calorimeter. The readout electronics for those sub-detectors remain inside their respective detector huts within the experimental Hall. All SHMS calorimeter PMT signals are fed into F250 FADCs configured to provide timing, integrated energy, pulse amplitude, and (optionally) pulse profile data as desired. The wire-chamber timing signals are digitized using multi-hit CAEN v1190 modules [CAEN1190].
@@ 5.1.1 JLab F250 Flash ADCs
The JLab F250 flash ADC modules are an FPGA-based design developed by the Jefferson Lab Fast Electronics group [JLABF250] and are used lab wide. Each F250 module provides 16 independent 50 input channels. The voltage at each input channel is continuously digitized into an 8 s ring buffer at 250 MHz, with a resolution of 12 bits, and a hardware adjustable full-scale range. When a module receives a readout trigger, digitized sample data stored in the ring buffer are processed in parallel without incurring front-end deadtime. In typical operation each ‘hit’ over a pre-programmed threshold is assigned an interpolated leading-edge threshold time (1 ns resolution), integrated energy (analogous to a charge-integrating ADC value), a peak amplitude, and a measurement of the DC offset (pedestal) present on the channel prior to the detected pulse. Full pulse-profile data for each hit may also be stored if desired. However, that mode increases the data rate by several orders of magnitude, and is generally used only for debugging or limited duration pulse characterization runs.
@@ 5.1.2 SHMS Triggers
An optional SHMS aerogel Cherenkov detector (AER), as detailed in Sec. 4.6, may also be installed. It employs seven PMTs on each side of its diffusion box. The signals from all 14 PMTs are handled analogously to the gas Cherenkov detectors, with each analog signal being split and read out by an individual FADC channel, with the second outputs being summed and discriminated to form an associated aerogel pre-trigger. The pre-trigger is routed to a scaler and v1190 TDC as well. A block diagram for the Cherenkov pre-triggers is presented in Fig. 51.
For the SHMS Preshower, described in Sec. 4.7, a pre-trigger is formed using the 28 analog signals from PMTs which are are split and summed in 3 groups of 4 rows, and 1 group of 2 rows. Each of the 4 group sums is read out by an FADC channel for cross checks. The 4 group sums are summed in turn to provide a total Preshower sum which is then discriminated and provides the SHMS PSh pre-trigger. Provision is made to generate independent pre-triggers for both low and high energy depositions in the Preshower layer (PSh_Lo and PSh_Hi, respectively) as seen in Fig. 52.
@@ 5.3 Auxiliary Data Collection
The standard method for slow controls data logging is through the Experimental Physics and Industrial Control System (EPICS) [EPICS]. EPICS is a system of open source software tools and applications used to provide control user interfaces and data logging for systems such as high- and low-voltage detector power supplies, target systems, spectrometer magnets, vacuum, and cryogenic systems, etc.
@@ 5.4 Online Hall C Computing Environment
All systems have direct access to the JLab centrally managed Scientific Computing resources. This includes multi-petabyte tape storage and online disk facilities, as well as a several thousand core compute farm for simulation and offline data analysis [JLABFARM].
@@ 6 Software
Hall C Data is analyzed by the Hall C analysis package hcana. This package does full event reconstruction for the SHMS used alone or in coincidence with other detectors. hcana is based on the modular Hall A analyzer [HallAAnalyzer] ROOT [ROOT] based C++ analysis framework. This framework provides for run time user configuration of histograms, ROOT tree contents, cuts, parameters and detector layout.
@@ 6.1 Online Monitoring
After each data taking run is started, a subset of the data is analyzed with hcana. An easily configurable histogram display GUI is used to view diagnostic histograms and compare them to reference histograms. The EPICS [EPICS] control system alarm handler is used to monitor experiment settings and beam conditions. This includes spectrometer magnet settings, detector high voltages, drift chamber gas, cryogenic systems and spectrometer vacuum.
@@ 7.3.1 Hodoscope Performance
Beam data confirmed the expectation that the detection efficiency for low momentum protons, for example, will be smaller than that for pions or electrons simply due to the reduced number of Cherenkov photons that particles close to their firing threshold will produce. This is exemplified by Figs. 62, 63 and 64.
@@ 7.3.2 Drift Chamber Performance
The SHMS drift chambers have proven to provide reliable tracking for electrons and hadrons across a broad range of momenta. The position resolution depends on how well the wire planes are aligned in the reconstruction software, but per-plane resolutions of 175 m are typical. The drift chambers have also performed very well at high rate, with tracking efficiencies exceeding 96%, even at pre-trigger rates over 2 MHz. The tracking efficiency as a function of the S1X hodoscope trigger rate (a good proxy for the overall event rate) in the SHMS can be seen in Figs. 65 and 66.
@@ 7.3.3 HGC Performance
Lastly, measurements of the efficiency across a variety of momentum settings can be used to verify the index of refraction of the Cherenkov media. The relationship between efficiency and momentum is fit with the equation [BrashAerogel]
where is the detector efficiency, is the momentum of the , and and are free parameters. Data taken in the range of 2.53 GeV/c to 5.05 GeV/c with the HGC filled with C4F10 yields an index of refraction of . This is in agreement with the accepted value of [C4F10Index]. Additional performance details are given in [RyanThesis].
@@ 8 Conclusion
Numerous experiments have completed and published high profile results since the SHMS was commissioned in 2017. This includes many high profile results on color transparency [jr:DBhetuwal_PRL2020, jr:DBhetuwal_PRC2023], the EMC effect [jr:AKarki_PRC2023], deuteron structure [jr:CYero_PRL2020] and proton structure [jr:RLi_Nature2022]. This also extends to detailed studies of the proton’s gravitational form factors [jr:BDuran_Nature2023]. Results on the F2 ratio for D/H[biswas2024newmeasurementsdeuteronproton] and the flavor dependence in charged pion SIDIS[bhatt2024flavordependencechargedpion] were recently submitted for publication.
Many more high profile scientific results are expected in the near future, with several experimental campaigns now completed and data analysis in advanced stages. Due to the design parameters of the SHMS, it could also be utilised extensively in an upgraded, 20+ GeV Jefferson Lab scenario. Some possible experiments that utilise the SHMS in a 20+ GeV era are outlined in the 22 GeV white paper [PrePrint:JLab22GeV].